OpenStudy (anonymous):

Solve for x: −3|x + 7| = −12

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

start by dividing both sides by -3

OpenStudy (anonymous):

x+4 = =12?

Vocaloid (vocaloid):

nope, not quite -12/(-3) = ?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

This is what i was given x = 5 \3, x = −19\3 x = −3, x = −11 x = −3, x = 11 No solution

Vocaloid (vocaloid):

good, that gives us |x + 7| = 4 now solve: x + 7 = 4 AND x + 7 = -4 you will get two x-values.

OpenStudy (anonymous):

x = −3, x = 11?

Vocaloid (vocaloid):

-3 is right 11 is not

Vocaloid (vocaloid):

x + 7 = -4 x = ?

OpenStudy (anonymous):

Would it be -11?

Vocaloid (vocaloid):

yup, so our answers are x = -3 and x = -11

OpenStudy (anonymous):

I get how the one is -3 but don't understand how it is -11. Do you have time and mind explaining?

Vocaloid (vocaloid):

sure x + 7 = -4 now we subtract 7 from each side x + 7 - 7 = -4 - 7 x = -4 - 7 = -11

OpenStudy (anonymous):

Cause when i did it i got 1.5714 that's why im a bit confused

OpenStudy (anonymous):

Ohhh! OK thank you so much for the help!